# MultiKernelBench: A Multi-Platform Benchmark for Kernel Generation

### v1

@@ 4.1. Methodology
DeepSeek-V3-0324 74.0 20.4 6.0 10.2 2.5 1.1 70.9 4.2 1.4 77
Qwen3-235B 66.3 17.9 5.3 0.7 0.7 0.4 89.8 3.85 2.5 64
Qwen2.5-Coder-32B 66.0 13.0 3.2 2.8 1.8 0.7 94.7 4.6 1.8 55
GPT-4o 97.5 19.6 3.9 2.5 1.8 0.4 97.2 4.9 4.6 75
Claude-Sonnet-4 92.3 28.1 8.8 5.3 2.1 0.4 99.6 5.9 4.9 103
DeepSeek-R1-0528 75.4 25.6 9.8 1.4 1.4 0.0 81.4 1.8 0.7 82
Qwen3-235B (think) 77.9 23.5 8.4 0.7 0.7 0.0 93.0 5.6 5.3 85
@@
DeepSeek-V3-0324 98.2 27.7 9.8 7.7 2.8 1.4 97.9 5.9 5.6 104
Qwen3-235B 97.2 25.6 7.7 3.9 2.1 1.4 100.0 6.3 4.2 97
Qwen2.5-Coder-32B 94.0 18.6 5.3 4.2 2.5 1.1 100.0 5.3 3.9 75
GPT-4o 99.6 26.7 9.8 3.2 2.1 0.7 100.0 4.6 4.6 95
Claude-Sonnet-4 97.9 31.2 11.2 8.1 3.9 2.1 100.0 5.9 5.6 117
@@
To ensure consistent and reproducible performance measurements, we restrict our evaluation of baseline implementations to PyTorch’s eager mode, excluding TorchDynamo and other compilation modes due to their platform-dependent behavior, which could introduce measurement inconsistencies. We report all metrics (Pass@k, Compilation@k, and $SpeedUp_{\alpha}@k$) for $k=1$ and $k=5$. For $k=1$, we use greedy decoding (temperature $=0.0$) to evaluate the model’s top-predicted solution. For $k=5$, we employ stochastic sampling with temperature $=0.2$ and top-$p=0.95$ to balance diversity and quality in the generated outputs, following established practices in prior code generation benchmarks (Chen et al., 2021; Cao et al., 2024). Following prior work (Ouyang et al., 2025), we do not apply stochastic sampling to reasoning models, as they entail significantly higher inference time and computational cost.
@@ 4.2. RQ1: Performance in Default Setting
Table 4(b) presents the overall performance of the selected LLMs across 285 kernel tasks under the default setting, where each task uses the add task as a one-shot example. For each platform, we report three key metrics—Compilation@k, Pass@k, and $SpeedUp_{1}@k$—for all selected models. Additionally, we include the total number of passed kernels for each model to provide a holistic view of overall performance. The best results for each metric are highlighted in bold to indicate the top-performing models. Overall, MultiKernelBench proves highly challenging: even the best-performing model, Claude-Sonnet-4, successfully completes only a subset of tasks (103 out of 855) under greedy decoding. While sampling search improves performance, a majority of tasks remain unsolved, underscoring the benchmark’s difficulty.
@@
Categories Claude-Sonnet-4 (%) DeepSeek-R1 (%) Qwen3-235B (think) (%) Avg Pass@1 (%)
@@
Activation 100.0 100.0 26.7 86.7 80.0 26.7 93.3 86.7 26.7 88.9
Broadcast 100.0 90.0 0.0 90.0 80.0 20.0 100.0 80.0 0.0 83.3
Convolution 100.0 0.0 0.0 58.8 0.0 0.0 52.9 0.0 0.0 0.0
Full Architecture 94.0 0.0 0.0 64.0 0.0 0.0 70.0 0.0 0.0 0.0
Fusion 85.0 4.0 3.0 78.0 6.0 5.0 81.0 4.0 2.0 4.7
Loss 85.7 42.9 28.6 100.0 42.9 42.9 100.0 57.1 42.9 47.6
Math 100.0 66.7 0.0 50.0 50.0 16.7 66.7 33.3 0.0 50.0
Matrix Multiply 100.0 70.6 5.9 88.2 82.4 5.9 94.1 76.5 5.9 76.5
Normalization 87.5 87.5 25.0 75.0 75.0 37.5 100.0 62.5 50.0 75.0
Optimizer 100.0 100.0 80.0 60.0 40.0 40.0 100.0 60.0 60.0 66.7
Pooling 100.0 66.7 50.0 83.3 33.3 0.0 66.7 16.7 16.7 38.9
Index 91.7 75.0 33.3 83.3 58.3 25.0 58.3 41.7 25.0 58.3
Resize 100.0 40.0 20.0 90.0 50.0 30.0 80.0 50.0 20.0 46.7
Reduce 80.0 80.0 0.0 100.0 100.0 20.0 100.0 80.0 20.0 86.7
@@
Platform Comparison. Model performance varies notably across platforms, primarily reflecting differences in training corpus coverage. On CUDA, GPT-4o achieves the highest Compilation@1 rate (97.5%), indicating strong syntactic alignment with CUDA conventions, while Claude-Sonnet-4 delivers the best Pass@1, Pass@5, and $SpeedUp_{1}@5$, suggesting superior optimization capabilities for CUDA code. In contrast, performance drops sharply on AscendC and Pallas for most models. DeepSeek-V3 shows the best AscendC result with a modest Pass@1 of 2.5%, while all other models remain below 2.1%. On the Pallas platform, Claude-Sonnet-4 leads with the highest Pass@1 (5.9%), whereas DeepSeek-R1, despite its strong CUDA performance, ranks lowest with only 1.8% Pass@1. These platform-specific discrepancies indicate that model effectiveness is highly correlated with the presence of relevant examples in the training data; the weaker performance on AscendC and Pallas likely stems from their limited representation in the pretraining corpora of current LLMs.
@@
Failure Analysis. On the CUDA platform, compilation errors are relatively uncommon, likely because CUDA is widely documented and has extensive coverage on the internet, despite its C-like syntax and complexity. However, generating functionally correct kernels remains a significant challenge for LLMs. Even Claude-Sonnet-4, the best-performing model on CUDA tasks, achieves only a 28.1% Pass@1 and 31.2% Pass@5 rate—most tasks still fail. We analyze the failure modes of compiled kernels and find that output mismatch is the dominant runtime error, accounting for 62.5% of failures. Output shape mismatches contribute 13.1%, while CUDA runtime errors account for 12.6%. These results highlight the difficulty LLMs face in reasoning about precise output semantics and data structure alignment, even in well-documented domains like CUDA.
@@
Activation 100.0 66.7 46.7 100.0 66.7 66.7 93.3 53.3 46.7 62.2
Broadcast 100.0 40.0 40.0 90.0 30.0 30.0 100.0 40.0 40.0 36.7
Fusion 100.0 1.0 1.0 91.0 2.0 1.0 96.0 0.0 0.0 1.0
@@ 4.3. RQ2: Category-wise Analysis
Category Comparison. For CUDA kernels, the top three models achieve the highest Pass@1 rates in the Activation, Reduce, and Broadcast categories, reaching 88.9%, 86.7%, and 83.3%, respectively. However, they fail entirely on the Convolution and Full Architecture categories, underscoring the complexity of these tasks. Notably, prior work (Ouyang et al., 2025) classifies Convolution and Activation at the same difficulty level, yet their actual Pass@1 rates differ sharply (88.9% vs. 0%), highlighting the need for our proposed fine-grained categorization. Fusion and Pooling also show low Pass@1 rates, at 4.7% and 38.9%, respectively. Overall, this category-wise analysis highlights substantial performance variation across different task types.

### v2

@@ 4.1. Methodology
DeepSeek-V3-0324 74.7 21.4 6.3 10.2 2.5 1.1 70.9 4.6 4.2 81
Qwen3-235B 66.3 22.5 5.3 0.7 0.7 0.4 89.8 3.9 1.8 77
Qwen2.5-Coder-32B 66.0 15.8 3.5 2.8 1.8 0.7 94.7 4.6 3.2 63
GPT-4o 97.5 23.2 5.3 2.5 1.8 0.4 97.2 6.3 2.8 89
Claude-Sonnet-4 92.3 47.0 20.4 5.3 2.1 0.4 99.6 8.4 7.7 164
DeepSeek-R1-0528 75.4 52.6 26.0 1.4 1.4 0.0 81.4 2.8 1.4 162
Qwen3-235B (think) 79.3 44.2 19.3 0.7 0.7 0.0 93.0 7.4 7.0 149
@@
DeepSeek-V3-0324 98.2 35.1 13.7 7.7 2.8 1.4 97.9 6.0 5.6 125
Qwen3-235B 97.2 43.9 13.0 3.9 2.1 1.4 100.0 6.3 4.2 149
Qwen2.5-Coder-32B 94.7 31.9 11.9 4.2 2.5 1.1 100.0 5.3 3.9 113
GPT-4o 99.6 32.3 12.6 3.2 2.1 0.7 100.0 9.8 8.8 126
Claude-Sonnet-4 97.9 55.8 26.0 8.1 3.9 2.1 100.0 10.5 8.8 200
@@
[…]
@@ 4.2. RQ1: Performance in Default Setting
Table 4(b) presents the overall performance of the selected LLMs across 285 kernel tasks under the default setting, where each task uses the add task as a one-shot example. For each platform, we report three key metrics—Compilation@k, Pass@k, and $SpeedUp_{1}@k$—for all selected models. Additionally, we include the total number of passed kernels for each model to provide a holistic view of overall performance. The best results for each metric are highlighted in bold to indicate the top-performing models. Overall, MultiKernelBench proves highly challenging: even the best-performing model, Claude-Sonnet-4, successfully completes only a subset of tasks (164 out of 855) under greedy decoding. While sampling search improves performance, a majority of tasks remain unsolved, underscoring the benchmark’s difficulty.
@@
Categories DeepSeek-R1 (%) Claude-Sonnet-4 (%) Qwen3-235B (think) (%) Avg Pass@1 (%)
@@
Activation 86.7 80.0 20.0 100.0 100.0 20.0 93.3 86.7 26.7 88.9
Broadcast 90.0 80.0 20.0 100.0 90.0 0.0 100.0 80.0 0.0 83.3
Convolution 58.8 32.4 5.9 100.0 0.0 0.0 58.8 8.8 0.0 13.7
Full Architecture 64.0 26.0 16.0 94.0 16.0 0.0 74.0 22.0 4.0 21.3
Fusion 78.0 59.0 44.0 85.0 50.0 37.0 81.0 49.0 32.0 52.7
Loss 100.0 42.9 28.6 85.7 42.9 14.3 100.0 57.1 42.9 47.6
Math 50.0 50.0 16.7 100.0 66.7 0.0 66.7 33.3 0.0 50.0
Matrix Multiply 88.2 76.5 5.9 100.0 70.6 5.9 94.1 76.5 5.9 74.5
Normalization 75.0 75.0 37.5 87.5 87.5 37.5 100.0 62.5 50.0 75.0
Optimizer 60.0 40.0 40.0 100.0 100.0 80.0 100.0 60.0 60.0 66.7
Pooling 83.3 33.3 0.0 100.0 66.7 50.0 66.7 16.7 0.0 38.9
Index 83.3 66.7 16.7 91.7 75.0 33.3 58.3 41.7 25.0 61.1
Resize 90.0 50.0 30.0 100.0 40.0 20.0 80.0 50.0 20.0 46.7
Reduce 100.0 100.0 20.0 80.0 80.0 0.0 100.0 80.0 20.0 86.7
@@
Platform Comparison. Model performance varies notably across platforms, primarily reflecting differences in training corpus coverage. On CUDA, GPT-4o achieves the highest Compilation@1 rate (97.5%), indicating strong syntactic alignment with CUDA conventions, while DeepSeek-R1 delivers the best Pass@1 and $SpeedUp_{1}@1$, suggesting superior optimization capabilities for CUDA code. In contrast, performance drops sharply on AscendC and Pallas for most models. DeepSeek-V3 shows the best AscendC result with a modest Pass@1 of 2.5%, while all other models remain below 2.1%. On the Pallas platform, Claude-Sonnet-4 leads with the highest Pass@1 (8.4%) and Pass@5 (10.5%), whereas DeepSeek-R1, despite its strong CUDA performance, ranks lowest with only 2.8% Pass@1. These platform-specific discrepancies indicate that model effectiveness is highly correlated with the presence of relevant examples in the training data; the weaker performance on AscendC and Pallas likely stems from their limited representation in the pretraining corpora of current LLMs.
@@
Failure Analysis. On the CUDA platform, compilation errors are relatively uncommon, likely because CUDA is widely documented and has extensive coverage on the internet, despite its C-like syntax and complexity. However, generating functionally correct kernels remains a significant challenge for LLMs. Even Claude-Sonnet-4, the best-performing model, achieves only a 47.0% Pass@1 and 55.8% Pass@5—indicating that many tasks still result in failure. We analyze the failure modes of compiled kernels and find that output mismatch is the dominant runtime error, accounting for 53.0% of failures. Output shape mismatches contribute 16.1%, while CUDA runtime errors account for 15.9%. These results highlight the difficulty LLMs face in reasoning about precise output semantics and data structure alignment, even in well-documented domains like CUDA.
@@
Activation 100.0 66.7 60.0 100.0 66.7 66.7 93.3 53.3 26.7 62.2
Broadcast 100.0 40.0 40.0 90.0 30.0 30.0 100.0 40.0 20.0 36.7
Full Architecture 98.0 4.0 4.0 92.0 4.0 4.0 98.0 6.0 2.0 4.7
Fusion 100.0 6.0 5.0 91.0 5.0 4.0 96.0 1.0 0.0 4.0
@@ 4.3. RQ2: Category-wise Analysis
Category Comparison. For CUDA kernels, the top three models achieve the highest Pass@1 rates in the Activation, Reduce, and Broadcast categories, reaching 88.9%, 86.7%, and 83.3%, respectively. In contrast, the Convolution and Full Architecture categories exhibit significantly lower pass rates—13.7% and 21.3%, respectively—highlighting the increased complexity of these tasks. Notably, prior work (Ouyang et al., 2025) classifies Convolution and Activation as having comparable difficulty, yet their actual Pass@1 rates diverge markedly (13.7% vs. 88.9%), reinforcing the need for our proposed fine-grained categorization. Pooling tasks also show relatively low performance, with a Pass@1 rate of 38.9%. Overall, this category-wise analysis reveals substantial variation in model performance across different task types.